OpenStudy (anonymous):

Integrate: ∫(cos(t))^2dt Integrate: ∫(cos(t))^2dt @Mathematics

OpenStudy (jamesj):

\( \cos 2x = \cos^2 x - \sin^2 x\) and therefore \( \cos 2x = 2 \cos^2 x - 1 \) Now you can rewrite the integrand in terms of functions you can integrate.

OpenStudy (anonymous):

I was trying to use reverse chain rule and integrate (u)^2 then integrate the inside cosine

OpenStudy (jamesj):

Well, you could use integration by parts twice. But doing what I'm suggesting is a bit shorter.

OpenStudy (anonymous):

Ah yea, ok

OpenStudy (anonymous):

Thanks